OpenStudy (swissgirl):

Prove that the function f:[1, infinity] --> (0, infinity) defined by f(x) =1/x maps onto (0, infinity)

OpenStudy (swissgirl):

What is wrong with this proof?

OpenStudy (swissgirl):

Suppose \( w \in (0, \inf) \). Choose \( x= \frac{1} {x} \). Then \( f(x) = \frac{1} {\frac{1} {w}} =w \) Therefore the function f is onto (0, infinity)

OpenStudy (anonymous):

sorry i cant help :( ask telliott99

OpenStudy (swissgirl):

hahaha alright @telliott99 can u help me?

OpenStudy (anonymous):

lol

OpenStudy (swissgirl):

Maybe all the proof is missing is that \(w \neq 0 \)

OpenStudy (swissgirl):

waittt a sec I messed up \( x= \frac{1} {w} \)

OpenStudy (zzr0ck3r):

that is not missing, you say w is in (0,inf) thus not 0

OpenStudy (swissgirl):

so what can be missing?

OpenStudy (zzr0ck3r):

o say that x is in the domain

OpenStudy (zzr0ck3r):

hmm if w is in (0,inf) 1/w is not always in (1,infinity)

OpenStudy (swissgirl):

ohhh so if w is a subset of (0. inf) we still must show that 1/w is a subset of (1 infinity) ok